OpenStudy (anonymous):

Which complex number has an absolute value of 5? A.–3 + 4i B.2 + 3i C.7 – 2i D.9 + 4i

OpenStudy (helder_edwin):

if z=a+ib then \[\large |z|=\sqrt{a^2+b^2} \]

OpenStudy (anonymous):

is it D ?

OpenStudy (helder_edwin):

use the formula. it should take u less than a couple of minutes

OpenStudy (helder_edwin):

if it were D then \[\large |9+4i|=\sqrt{9^2+4^2}=\sqrt{81+16}=\sqrt{97}=5?\]

OpenStudy (anonymous):

oh b

OpenStudy (helder_edwin):

do the math. do u get 5?

OpenStudy (helder_edwin):

use the formula. did u see what i did? we r talking about the absolute value of a complex number